# AttnLRP: Attention-Aware Layer-wise Relevance Propagation for Transformers

### v1

@@ A.3 Proofs
In the following, we provide proofs for the rules presented in the main paper.
@@
$R_{i\leftarrow j}=\begin{cases}(x_{i}-s_{i}x_{i})\ R^{l+1}_{i}&\text{for }i=j\\ -s_{i}x_{i}\ R^{l+1}_{j}&\text{for }i\neq j\end{cases}$
@@
$R^{l}_{i}=\sum_{j}R_{i\leftarrow j}=x_{i}(R^{l+1}_{i}-s_{i}\sum_{j}R^{l+1}_{j})$
@@
$\frac{\partial f_{j}}{\partial x_{i}}=\prod_{k\neq i}^{N}x_{k}$
@@
$f_{j}(\textbf{x})=\sum_{i}^{N}\frac{\partial f_{j}}{\partial x_{i}}x_{i}+\tilde{b}_{j}=N\prod_{k}^{N}x_{k}+\tilde{b}_{j}=Nf_{j}(\textbf{x})+\tilde{b}_{j}$
@@
$R_{i\leftarrow j}=\left(f_{j}+\frac{\tilde{b}_{j}}{N}\right)\frac{R_{j}}{Nf_{j}(\textbf{x})+\tilde{b}_{j}}=\frac{1}{N}R_{j}$
@@
Consider the equation for matrix multiplication, where we treat the terms as single input variables:
$\textbf{O}_{jp}=\sum_{i}\left(\textbf{A}_{ji}\textbf{V}_{ip}\right)$
@@
$R_{jip}(\textbf{A}_{ji},\textbf{V}_{ip})=\textbf{A}_{ji}\textbf{V}_{ip}\frac{R_{jp}}{\textbf{O}_{jp}}$
Next, we decompose the individual terms using the derived rule from the previous Section A.3.2
$R_{ji}=\sum_{p}R_{ji\leftarrow jip}=\sum_{p}\frac{1}{2}R_{jip}=\sum_{p}\textbf{A}_{ji}\textbf{V}_{ip}\frac{R_{jp}}{2\ \textbf{O}_{jp}}$
The proof for $\textbf{V}_{ip}$ follows a similar approach.
@@
$f_{j}(\textbf{x})=\frac{\partial f_{i}}{\partial x_{i}}\Bigr{|}_{\tilde{x}_{i}=0}x_{i}+\tilde{b_{j}}=\frac{x_{i}}{\varepsilon}+\tilde{b_{j}}$
To enforce a strict notion of the conservation property (3), the bias term $\tilde{b_{j}}$ can be excluded or evenly distributed across the input variables. Because we have only a single input variable, the bias can be considered as part of $x_{i}$.
$R_{i}=\left(\frac{x_{i}}{\varepsilon}+\tilde{b_{j}}\right)\frac{R_{j}}{\frac{x_{i}}{\varepsilon}+\tilde{b_{j}}}$ (28)
@@ B.1 Models and Datasets
For ImageNet faithfulness, we utilized the pretrained Vision Transformer 16-B weights of the PyTorch model zoo Paszke et al., (2019). We randomly selected 3200 samples such that the standard error of mean converges to below 1% of the mean value.
For Wikipedia, SQuAD v2 and IMDB faithfulness, we evaluated the pretrained Llama 2-7b and Flan-T5 weights hosted on huggingface Wolf et al., (2019) on 4000 randomly selected validation dataset samples (fixed set for all baselines). Both models are casted to the Brain Floating Point (BFLOAT16) half-precision format to save memory consumption and to simulate real-world scenarios. Further, for Wikipedia next word prediction we evaluated the model on a context size of 512 (from beginning of article until context length is reached), while the context size in SQuAD v2 varies between 169 to 4060. Although Flan-T5 was trained on a smaller context size of 2000 tokens, the relative positional encoding allows it to handle longer context sizes with at least 8192 tokens Shaham et al., (2023). Additionally for SQuAD v2 faithfulness and IoU, we utilized the following prompt:
@@
In order to prevent out-of-distribution samples, we only perturb the tokens inside the context [text of dataset sample]. Since the length of the SQuAD v2 questions varies and the length of the validation dataset is small, the standard error of mean is higher in Table 1.
For IMDB, we added a last linear layer to Llama 2-7b and finetuned the model, which achieves 93% accuracy on the validation dataset.
@@ B.2 Input Perturbation Metrics
Results of these processes are perturbed input sets of $\mathcal{X}^{F}_{MoRF}=(\textbf{x}_{MoRF}^{0},\textbf{x}_{MoRF}^{1},...,\textbf{x}_{MoRF}^{N-1})$ and $\mathcal{X}^{F}_{LeRF}=(\textbf{x}_{LeRF}^{0},\textbf{x}_{LeRF}^{1},...,\textbf{x}_{LeRF}^{N-1})$. It is notable that by using $\textbf{g}^{I}$, $\mathcal{X}^{I}_{LeRF}=\mathcal{X}^{F}_{MoRF}$ and vice versa. By feeding these sets to the model and computing the corresponding logit output $f_{j}$, a curve will be induced and consequently the area under the curve can be calculated:
@@
By using $\mathcal{X}^{I}_{MoRF}$, $A^{I}_{MoRF}=A^{F}_{LeRF}$ can be computed similarly. A faithful explainer results in a low value of $A^{F}_{MoRF}$ or $A^{I}_{LeRF}$. Further, a faithful explainer is expected to have large $A^{F}_{LeRF}$ or $A^{I}_{MoRF}$ values.
@@
Figure 4: Comparison of the AttnLRP (ours) with the $\gamma$-rule, Grad×$\times$×AttnRoll Chefer et al., 2021a , AtMan Deb et al., (2023), and SmoothGrad Smilkov et al., (2017) techniques through the perturbation experiment (faithfulness) on the ViT-B-16 using 3200 random samples of ImageNet. From left to right, the plots correspond to $f_{j}(\mathcal{X}^{F}_{LeRF})-f_{j}(\mathcal{X}^{F}_{MoRF})$ (large area is good), $f_{j}(\mathcal{X}^{F}_{MoRF})$ (steep decline is good), and $f_{j}(\mathcal{X}^{F}_{LeRF})$ (slow decline is good). “AUC” denotes the Area under Curve.
@@ B.3 Hyperparameter search for Baselines
B.4 LRP Composites for ViT
Applying the $\varepsilon$-rule on all linear layers inside LLMs is sufficient to obtain faithful and noise-free attributions. However, for the Vision Transformers, we apply the $\gamma$-rule on all linear layers (including the convolutional layers) outside the attention module. Since the $\gamma$-rule has a hyperparameter, the work Pahde et al., (2023) proposed to tune the parameter using a grid-search. This optimization search (or in LRP known as composite search) is computational highly demanding.
The Vision Transformer consists of many linear layers. Our proposed approach is to use different $\gamma$ values across different layer types. According to Vaswani et al., (2017) the attention module consists of several linear layers which we refer to as LinearInputProjection:
@@ B.4 LRP Composites for ViT
The most faithful composite, that we obtained for AttnLRP and CP-LRP, is in Table 4. More details over the statistics of the conducted experiments are available in Figures 12, 13, 14, 15, 16.
Table 4: Proposed composite for the AttnLRP and CP-LRP methods used for the Vision Transformer.
@@
Figure 5: Comparison of four different LRP variants computed on a Llama 2-7b model. The given section is from the Wikipedia article on Mount Everest. The model is expected to provide the next answer token for the question ‘How high did they climb in 1922? According to the text, the 1922 expedition reached 8,’. For the correctly predicted token 3 is the attribution computed. Distributing the bias uniformely on the input variables (Softmax Distribute Bias) or applying the identity rule (Softmax Identity Rule) leads to numerical instabilities. AttnLRP highlights the correct token the strongest, while CP-LRP focuses strongly on the start-of-sequence <s> token and exhibits more background noise e.g. irrelevant tokens such as ‘Context’, ‘attracts’, ‘Everest’ are highlighted, while AttnLRP does not highlight them or assigns negative relevance.
@@
Figure 6: Explanation heatmaps of the methods used for the perturbation experiments on the Vision Transformer. A checkerboard effect is visible for almost every method, especially in AttnRoll Abnar and Zuidema, (2020), G×$\times$×AttnRoll Chefer et al., 2021a , and AtMan Deb et al., (2023). We improve upon CP-LRP Ali et al., (2022) by applying the $\gamma$-rule as described in B.4. While qualitatively CP-LRP (with $\gamma$ extension for ViT) and AttnLRP give similar explanations, quantitative results in Table 1 show a consistent improvement of AttnLRP over CP-LRP in terms of faithfulness. Moreover, the detailed investigation of the processes within the attention model can be investigated with AttnLRP only, while it is not possible with CP-LRP. We leave these further explorations for future work. The quantitative experiments (faithfulness test) validate this observation, see Table 5.
B.5 Attributions on SQuAD v2
In Figure 7, we illustrate attributions for different state-of-the-art methods computed for the first token of the answer (highlighted in red). The Flan-T5 model correctly predicts 182 million, where the first explained token is 18.

### v2

@@ A.3 Proofs
In the following, we provide proofs for the rules presented in the main paper, and that the application of the $\varepsilon$-rule on bi-linear matrix multiplication violates the conservation property.
@@
$R_{i\leftarrow j}^{(l-1,l)}=\begin{cases}(x_{i}-s_{i}x_{i})\ R^{l}_{i}&\text{for }i=j\\ -s_{i}x_{i}\ R^{l}_{j}&\text{for }i\neq j\end{cases}$
@@
$R^{l-1}_{i}=\sum_{j}R_{i\leftarrow j}^{(l-1,l)}=x_{i}(R^{l}_{i}-s_{i}\sum_{j}R^{l}_{j})$
In Appendix A.2.4, we discuss the implications of vanishing gradients and temperature scaling on attributing the softmax function.
@@
$\frac{\partial f_{j}}{\partial x_{ji}}=\prod_{k\neq i}^{N}x_{jk}$
@@
$f_{j}(\textbf{x})=\sum_{i}^{N}\frac{\partial f_{j}}{\partial x_{ji}}x_{ji}+\tilde{b}_{j}=N\prod_{k}^{N}x_{jk}+\tilde{b}_{j}=Nf_{j}(\textbf{x})+\tilde{b}_{j}$
@@
$R_{ji\leftarrow j}^{(l-1,l)}=\left(f_{j}+\frac{\tilde{b}_{j}}{N}\right)\frac{R_{j}^{l}}{Nf(\textbf{x})_{j}+\tilde{b}_{j}}=\frac{1}{N}R_{j}^{l}$
Since each input with index $ji$ at layer $l-1$ is only connected to one output with index $j$ at layer $l$, we have only a single relevance propagation message. Hence, it follows from Equation (2):
$R_{ji}^{l-1}=R_{ji\leftarrow j}^{(l-1,l)}=\frac{1}{N}R_{j}^{l}$
@@
Consider the equation for matrix multiplication, where we treat the terms as single input variables by substituting them with $\textbf{u}_{jip}=\textbf{A}_{ji}\textbf{V}_{ip}$
$\textbf{O}_{jp}=\sum_{i}\textbf{A}_{ji}\textbf{V}_{ip}=\sum_{i}\textbf{u}_{jip}$
@@
$R_{jip\leftarrow jp}^{(l-1,l)}\propto\textbf{u}_{jip}$
This can also be seen, by noticing that $\sum_{i}\textbf{u}_{jip}$ is a linear operation, since a single variable is left. Hence, it can be regarded as a linear layer (6) characterized with constant weights of one and a bias of zero. We already derived the solution to applying a linearization (3.1.1) to a linear layer: the $\varepsilon$-rule (8). Therefore, the solution is:
$R_{jip\leftarrow jp}^{(l-1,l)}=\textbf{u}_{jip}\frac{R_{jp}^{l}}{\textbf{O}_{jp}+\varepsilon}=\textbf{A}_{ji}\textbf{V}_{ip}\frac{R_{jp}^{l}}{\textbf{O}_{jp}+\varepsilon}$
Since each input with index $jip$ at layer $l-1$ is only connected to one output with index $jp$ at layer $l$, we have only a single relevance propagation message. Hence, it follows from equation (2):
$R_{jip}^{l-1}=R_{jip\leftarrow jp}^{(l-1,l)}$
Next, we decompose the individual terms $\textbf{u}_{jip}$ using the uniform rule from the previous Section A.3.2 to obtain relevance messages for $\textbf{A}_{ji}$:
$R_{ji\leftarrow jip}^{(l-1,l-1)}=\frac{1}{2}R_{jip}^{l-1}$
Each input $\textbf{A}_{ji}$ is connected to $p$ outputs $\textbf{u}_{jip}$. Hence, to obtain the relevance values attributed to $\textbf{A}_{ji}$, we must aggregate all relevance messages from output ${jip}$ to inputs ${ji}$ via Equation (2):
$R_{ji}^{l-1}=\sum_{p}R_{ji\leftarrow jip}^{(l-1,l-1)}=\sum_{p}\frac{1}{2}R_{jip}^{l-1}$
$R_{ji}^{l-1}=\sum_{p}\textbf{A}_{ji}\textbf{V}_{ip}\frac{R_{jp}^{l}}{2(\textbf{O}_{jp}+\varepsilon)}$
Because $\varepsilon\ll|\textbf{O}_{jp}|$, we simplify the final solution:
$R_{ji}^{l-1}=\sum_{p}\textbf{A}_{ji}\textbf{V}_{ip}\frac{R_{jp}^{l}}{2\ \textbf{O}_{jp}+\varepsilon}$
The proof for $\textbf{V}_{ip}$ follows a similar approach by summing over the $j$ indices instead of $p$. In Appendix A.3.5, we proof that this rule does not violate the conservation property (3) in contrast to the standard $\varepsilon$-rule.
@@
$f_{i}(\textbf{x})=\frac{\partial f_{i}}{\partial x_{i}}\Bigr{|}_{\tilde{x}_{i}=0}x_{i}+\tilde{b}_{i}=\frac{x_{i}}{\varepsilon}+\tilde{b}_{i}$
To enforce a strict notion of the conservation property (3), the bias term $\tilde{b}_{i}$ can be excluded or evenly distributed across the input variables. Because we have only a single input variable, the bias can be considered as part of $x_{i}$.
$R_{i}^{l-1}=\left(\frac{x_{i}}{\varepsilon}+\tilde{b}_{i}\right)\frac{R_{i}^{l}}{\frac{x_{i}}{\varepsilon}+\tilde{b}_{i}}$ (29)
@@
A.3.5 Violation of the Conservation Property in bi-linear Matrix Multiplication
In the following we proof that the application of the $\varepsilon$-rule (8) without the uniform rule (14) on bi-linear matrix multiplication violates the conservation property (3). We reiterate and generalize the Lemma 3 of Chefer et al., 2021b which establishes that 00-LRP ($\varepsilon=0$) violates conservation.
Recall, that matrix multiplication is defined as:
$\textbf{O}_{jp}=\sum_{i}\textbf{A}_{ji}\textbf{V}_{ip}$
The $\varepsilon$-rule is the solution to applying a linearization (3.1.1) to a linear layer (6). For computing relevance values for $\textbf{A}_{ji}$ using the $\varepsilon$-rule, we treat $\textbf{V}_{ip}$ as a constant weight matrix with zero bias, and similarly for attributing $\textbf{A}_{ji}$. The derived relevance propagation rules are given by:
$\tilde{R}_{ji}^{l-1}(\textbf{A}_{ji})=\sum_{p}\textbf{A}_{ji}\textbf{V}_{ip}\frac{R_{jp}^{l}}{\textbf{O}_{jp}+\varepsilon}$
$\tilde{R}_{ip}^{l-1}(\textbf{V}_{ip})=\sum_{j}\textbf{A}_{ji}\textbf{V}_{ip}\frac{R_{jp}^{l}}{\textbf{O}_{jp}+\varepsilon}$
The conservation property (3) states, that the total relevance at layer $l$ must be equal to the total relevance at layer $l-1$.
The total relevance at layer $l$ is given by
$R^{l}=\sum_{j,p}R_{jp}^{l}$
and the total relevance at layer $l-1$ is computed by:
$R^{l-1}=\sum_{j,i}\tilde{R}_{ji}^{l-1}+\sum_{i,p}\tilde{R}_{ip}^{l-1}=2\sum_{j,p}\frac{\textbf{O}_{jp}}{\textbf{O}_{jp}+\varepsilon}R_{jp}^{l}\approx 2R^{l}$
with $\varepsilon\ll|\textbf{O}_{jp}|$. This results in a violation of the conservation property as $R^{l}\neq R^{l-1}$. However, by employing Proposition 3.3 (15), that is a sequential application of the $\varepsilon$-rule and uniform rule (14), we ensure conservation by dividing with the factor $2$:
$R^{l-1}=\sum_{j,i}R_{ji}^{l-1}+\sum_{i,p}R_{ip}^{l-1}=\sum_{j,p}\frac{\textbf{O}_{jp}}{\textbf{O}_{jp}+\varepsilon}R_{jp}^{l}\approx R^{l}$
It is evident that $\varepsilon$ absorbs a negligible proportion of the relevance to safeguard numerical stability. The proof is also valid for the $z^{+}$-rule (25), where only positive contributions are taken into consideration.
Figure A.4: Comparison of four different LRP variants computed on a LLaMa 2-7b model. The given section is from the Wikipedia article on Mount Everest. The model is expected to provide the next answer token for the question ‘How high did they climb in 1922? According to the text, the 1922 expedition reached 8,’. For the correctly predicted token 3 the attribution is computed. Distributing the bias uniformely on the input variables (Softmax Distribute Bias) or applying the identity rule (Softmax Identity Rule) leads to numerical instabilities. For “Softmax Distribute Bias” and “Softmax Identity Rule”, we applied AttnLRP rules on all layers except for the softmax function. AttnLRP highlights the correct token the strongest, while CP-LRP focuses strongly on the start-of-sequence <s> token and exhibits more background noise e.g. irrelevant tokens such as ‘Context’, ‘attracts’, ‘Everest’ are highlighted, while AttnLRP does not highlight them or assigns negative relevance.
@@ B.1 Models and Datasets
For ImageNet faithfulness, we utilized the pretrained Vision Transformer B-16, L-16 and L-32 weights of the PyTorch model zoo Paszke et al., (2019). We randomly selected 3200 samples (fixed set for all baselines) such that the standard error of mean converges to below 1% of the mean value.
For Wikipedia and IMDB faithfulness, we evaluated the pretrained LLaMa 2-7b hosted on huggingface Wolf et al., (2019) on 4000 randomly selected validation dataset samples (fixed set for all baselines). For SQuAD v2, we utilize the pretrained Flan-T5 and Mixtral 8x7b weights hosted on huggingface. Further, for Wikipedia next word prediction we evaluated the model on a context size of 512 (from beginning of article until context length is reached), while the context size in SQuAD v2 varies between 169 to 4060. Although Flan-T5 was trained on a smaller context size of 2000 tokens, the relative positional encoding allows it to handle longer context sizes with at least 8192 tokens Shaham et al., (2023).
All computations are performed in the Brain Floating Point (bfloat16) half-precision format to save memory consumption. bfloat16 trades precision for a higher dynamic range than standard float16, and hence prevents numerical errors due to overflow. In this regard, the impact of quantized number formats on (Attn)LRP attributions remains a topic to be investigated. In addition, all linear weights in Mixtral 8x7b are quantized to the 4 bit integer format using bitsandbytes Dettmers et al., (2024) (but computation still performed in bfloat16).
SQuAD v2 encompasses numerous questions that are either unanswerable or subject to incorrect predictions by the model. Consequently, only instances where the model accurately predicts the correct response are considered. Additionally, only the testset is utilized to mitigate a potential overfitting bias during the training phase, if applicable. Finally, for SQuAD v2 top-1 accuracy and IoU, we utilized the following prompt:
@@
Flan-T5 does not require a system prompt, while for Mixtral 8x7b we use before the context the system prompt:
Use the context to answer the question. Use few words.
Because Flan-T5 typically provides the correct answer directly, we explain the first token of the answer only. Conversely, Mixtral 8x7b generates full sentences; within these, we identify the positions of the answer tokens and explain all tokens that constitute the correct answer only. To achieve this, we calculate heatmaps for each answer token and add these heatmaps to produce the final heatmap. For gradient-based methods, this process can be parallelized by initiating the backward pass at the designated token positions with the logit output for LRP and with the value 1 for all other baselines, while initializing the remaining output tokens with zero.
For IMDB, we added a last linear layer to a frozen LLaMa 2-7b model and finetuned only the last layer, which achieves 93% accuracy on the validation dataset.
@@ B.2 Input Perturbation Metrics
Results of these processes are perturbed input sets of $\mathcal{X}^{F}_{MoRF}=(\textbf{x}_{MoRF}^{0},\textbf{x}_{MoRF}^{1},...,\textbf{x}_{MoRF}^{N-1})$ and $\mathcal{X}^{F}_{LeRF}=(\textbf{x}_{LeRF}^{0},\textbf{x}_{LeRF}^{1},...,\textbf{x}_{LeRF}^{N-1})$. By feeding these sets to the model and computing the corresponding logit output $f_{j}$, a curve will be induced and consequently the area $A$ under the curve can be calculated:
@@
It is notable that the area below the least relevant order insertion curves are identical to the most relevant order flipping curves and that the area below the least relevant order flipping curves are identical to the most relevant order insertion curves. Hence, by using $\mathcal{X}^{I}_{MoRF}$, $A^{I}_{MoRF}=A^{F}_{LeRF}$ can be computed similarly. A faithful explainer results in a low value of $A^{F}_{MoRF}$ or $A^{I}_{LeRF}$. Further, a faithful explainer is expected to have large $A^{F}_{LeRF}$ or $A^{I}_{MoRF}$ values.
@@
Figure B.5: Comparison of the AttnLRP (ours) with the $\gamma$-rule, Grad×$\times$×AttnRoll Chefer et al., 2021a , AtMan Deb et al., (2023), and SmoothGrad Smilkov et al., (2017) techniques through the perturbation experiment (faithfulness) on the ViT-B-16 using 3200 random samples of ImageNet. From left to right, the plots correspond to $f_{j}(\mathcal{X}^{F}_{LeRF})-f_{j}(\mathcal{X}^{F}_{MoRF})$ (large area is good), $f_{j}(\mathcal{X}^{F}_{MoRF})$ (steep decline is good), and $f_{j}(\mathcal{X}^{F}_{LeRF})$ (slow decline is good). “AUC” denotes the Area under Curve.
@@ B.3 Hyperparameter search for Baselines
For LLMs, we have not noticed a significant impact on the heatmaps for different discard threshold values of G×$\times$×AttnRoll. For AttnRoll, the impact is minimal. Hence, we choose the default value of 1 (nothing is discarded, as proposed in the original works Abnar and Zuidema, (2020); Chefer et al., 2021a ).
Regarding SQuAD v2, we set AtMan’s $p=0.7$ for Mixtral 8x7b and $p=0.9$ for Flan-T5-XL. For SmoothGrad, we set $\sigma=0.1$ for Mixtral 8x7b and Flan-T5-XL.
B.4 Impact of Model Architectural Choices on AttnLRP Performance
We evaluated AttnLRP on three model classes that incorporate different types of layers.
Flan-T5: This encoder-decoder architecture employs self-attention and cross-attention layers (33). The FFN layers are a sequential application of linear layers with GELU non-linearities inbetween.
LLaMa 2: This decoder architecture utilizes only self-attention layers (33). However, in the FFN layers, we have an additional element-wise non-linear weighting with a SiLU non-linearity (34).
Mixtral 8x7b: This mixture of experts model uses self-attention layers (33) and FFN layers with non-linear weighting (34) like the LLaMa 2. In addition, there are FFN routing layers with a softmax weighting (35).
[…]
where $\textbf{W}_{q}$, $\textbf{W}_{k}$, $\textbf{W}_{v}$, $\textbf{W}_{1}$, $\textbf{W}_{2}$, $\textbf{W}_{g}$ are linear weight parameters, ($\odot$) is element-wise multiplication, $i\in\mathbb{N}$ the number of expert FFN layers and TopK is returning the top-k elements.
In Table B.4, we study the impact of AttnLRP rules w.r.t. CP-LRP on all three different layer types (33), (34) and (35). We start as baseline with all rules of CP-LRP on all layer types, then we successively substitute CP-LRP rules with AttnLRP rules for specific layer types.
For CP-LRP, we use the rules described in Appendix A.2.2. In the original work of Ali et al., (2022), FFN layers weighted with non-linearities and routing layers are not discussed. Analogously to the argumentation in their work, we regard the non-linearity as constant weight and attribute only through the FFN path using the $\varepsilon$-rule.
As demonstrated in Table B.4, the application of AttnLRP rules enhances the performance across all layers, regardless of their type. Moreover, the rate of improvement increases with the number of non-linearities present in the model.
Table B.4: This Table is an extension of Table 1: Faithfulness scores as area between the least and most relevant order perturbation curves on LLaMa 2 alongside the top-1 accuracy and IoU in parenthesis for Flan-T5 and Mixtral 8x7b. We start as baseline with all rules of CP-LRP on all layer types, then we successively substitute CP-LRP rules with AttnLRP rules for specific layer types if they exist in the model. We observe that AttnLRP’s improvement is not confined to the attention mechnism alone, but all layers that contain operations that are not attributable with other LRP variants. The more complex the architecture, the better the performance of AttnLRP compared to CP-LRP.
Method on Layer LLaMa 2-7b Mixtral 8x7b Flan-T5-XL
IMDB $\uparrow$ Wikipedia $\uparrow$ SQuAD v2 $\uparrow$ SQuAD v2 $\uparrow$
Baseline (All Layers)
CP-LRP $1.72$ $7.85$ $0.50\ (0.40)$ $0.90\ (0.83)$
+ Attention Mechanism
AttnLRP $2.09$ $9.49$ $0.70\ (0.53)$ $0.94\ (0.84)$
+ FFN $\times$ Non-Linearity
AttnLRP $2.50$ $10.93$ $0.78\ (0.57)$
+ Routing Layer
AttnLRP - - $0.96\ (0.72)$
B.5 LRP Composites for ViT
Applying the $\varepsilon$-rule on all linear layers inside LLMs is sufficient to obtain faithful and noise-free attributions. However, for the vision transformers, we apply the $\gamma$-rule on all linear layers (including the convolutional layers) outside the attention module. Since the $\gamma$-rule has a hyperparameter, the work Pahde et al., (2023) proposed to tune the parameter using a grid-search. This optimization search (or in an LRP context known as composite search) is computational highly demanding.
The vision transformer consists of many linear layers. Our proposed approach is to use different $\gamma$ values across different layer types.
According to Vaswani et al., (2017) the attention module consists of several linear layers which we refer to as LinearInputProjection.
@@ B.5 LRP Composites for ViT
The most faithful composite, that we obtained for AttnLRP and CP-LRP, is in Table B.5. More details over the statistics of the conducted experiments are available in Figures B.14, B.15, B.16, B.17, B.18.
Table B.5: Proposed composite for the AttnLRP and CP-LRP methods used for the Vision Transformer.
@@
Figure B.6: Explanation heatmaps of the methods used for the perturbation experiments on the vision transformer ViT-B-16. A checkerboard effect is visible for almost every method, especially in AttnRoll Abnar and Zuidema, (2020), G×$\times$×AttnRoll Chefer et al., 2021a , and AtMan Deb et al., (2023). We improve upon CP-LRP Ali et al., (2022) by applying the $\gamma$-rule as described in B.5. While qualitatively $\gamma$CP-LRP (with $\gamma$ extension for ViT) and AttnLRP give similar explanations, quantitative results in Table 1 and Table B.6 show a consistent improvement of AttnLRP over $\gamma$CP-LRP in terms of faithfulness. Moreover, attributing query and key linear layers within the attention layer is possible with AttnLRP only, while it is not possible with CP-LRP. We leave these further explorations for future work.
B.6 Additional Perturbation Evaluations on Vision Transformers
Table B.6 presents additional perturbation results for the vision transformers ViT-L-16 and ViT-L-32 evaluated on ImageNet. Our method surpasses all comparative baselines, while the proposed enhancement, $\gamma$CP-LRP (which applies the $\gamma$-rule across all linear layers for CP-LRP Ali et al., (2022)), remains highly competitive. In more complex model architectures that incorporate a greater variety of non-linearities, our method demonstrates more superiority, as elaborated in Appendix B.4. Tuning the $\gamma$-parameter for $\gamma$CP-LRP and AttnLRP in a grid-search (see Appendix B.5) resulted for both models in the same composite described in Table B.5. However, there is no assurance that other models share the same $\gamma$ parameters.
Table B.6: Faithfulness scores as area between the least and most relevant order perturbation curves Blücher et al., (2024) for ViT-L-16 and ViT-L-32 on ImageNet. For ViT-L-16, we set for SmoothGrad $\sigma=0.01$, for AtMan $p=1.0$ and $t=0.1$, for AttnRoll $dt=0.90$, and for G×$\times$×AttnRoll $dt=0.92$. For ViT-L-32, we set for SmoothGrad $\sigma=0.01$, for AtMan $p=1.0$ and $t=0.1$, for AttnRoll $dt=0.95$, and for G×$\times$×AttnRoll $dt=1.0$.
Method ViT-L-16 ViT-L-32
ImageNet $\uparrow$ ImageNet $\uparrow$
Random --0.01± ${\,\pm\,0.01}$ 0.01 --0.01± ${\,\pm\,0.01}$ 0.01
[…]
AttnLRP (ours) --$\textbf{7.17}{\,\pm\,0.04}$-$\textbf{6.06}{\,\pm\,0.04}$
B.7 Attributions on SQuAD v2
In Figure B.7 and Figure B.8, we illustrate attributions on the Mixtral 8x7b for different state-of-the-art methods on the SQuAD v2 dataset. In Figure B.9, we depict attributions for Flan-T5-XL. For comparison, we also visualize a random attribution with Gaussian noise.
The similarity between AttnLRP and CP-LRP in Flan-T5-XL are in line with the quantitative evaluation from Table 1, which shows a small, but consistent advantage of AttnLRP over CP-LRP wrt. top-1 accuracy, while in Mixtral 8x7b and LLaMa 2, AttnLRP substantially outperforms, which is also visible in the heatmaps. This is due to the different number of non-linearities present in the models: Flan-T5-XL consists only of standard attention layers, while LLaMa 2 and Mixtral 8x7b have additional FFN layers with non-linear weighting or routing layers making the attribution process more difficult for CP-LRP. This effect is studied in Appendix B.4. In general, gradient-based methods such as G×$\times$×I, SmoothGrad, IG and Grad-CAM are noisy and often not informative. Attention Rollout and Grad×$\times$×Attention Rollout suffer from background noise. While the performance of AtMan is in some cases excellent as in Figure B.9, the method fails in other cases as in Figure B.7.
In Figure B.7 and B.8 for Mixtral 8x7b, most methods fail to highlight the correct answer tokens most strongly, except AttnLRP, confirming the quantitative evaluation from Table 1.